OpenStudy (faiqraees):

If \[\large\rm x=e^t\ \] Show that (Done) \[\large\rm x\frac{dy}{dx}=\frac{dy}{dt} \ \] And (Need help on this) \[\large\rm x\frac{d^2y}{dx^2}=\frac{d^2y}{dt^2}-\frac{dy}{dt} \] And please also explain if we are going to describe \[\large\rm \frac{d^2y}{dx^2},\frac{dy^2}{dx^2},d^2y \] in English, what would it be? (Like change in y with respect to.....)

OpenStudy (faiqraees):

@zepdrix

OpenStudy (loser66):

Is it not that \(\dfrac{d}{dt}(\dfrac{dy}{dt})=\dfrac{dy}{dx}+x\dfrac{d^2y}{dx^2}\) and you got the answer?

OpenStudy (faiqraees):

How do you got that. Can you explain?

OpenStudy (loser66):

\(\dfrac{d^2y}{dx^2}\) is just second derivative of y with respect to x.

OpenStudy (loser66):

The left hand side is just a factor of x and dy/dx. Take product rule.

OpenStudy (loser66):

@zepdrix now, your turn. hehehe...

zepdrix (zepdrix):

x=e^t where is y...?

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Loser66 \(\dfrac{d^2y}{dx^2}\) is just second derivative of y with respect to x. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah but what does it mean in English? Like dy/dx is said to be change in y with respect to change in x

OpenStudy (loser66):

@zepdrix Please, explain him.

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Loser66 Is it not that \(\dfrac{d}{dt}(\dfrac{dy}{dt})=\dfrac{dy}{dx}+x\dfrac{d^2y}{dx^2}\) and you got the answer? \(\color{#0cbb34}{\text{End of Quote}}\) It is acquired by doing right? \[\large\rm \frac{d}{dx} (x\frac{dy}{dx})=\frac{d}{dt} (\frac{dy}{dt} ) \]

OpenStudy (loser66):

yup

OpenStudy (loser66):

about second derivative!! I think you should understand what it means!!

OpenStudy (loser66):

Suppose you have a function y =x^2, then dy / dx is the tangent of the curve, right?

zepdrix (zepdrix):

derivative measures change in y with respect to x, so, second derivative measures change in (change in y with respect to x). It's measuring the change in the change. It makes a little more sense if you think of it in terms of movement I suppose. Like if you're measuring the movement of a particle over time... derivative measures the moment over time (we call this velocity), second derivative measures the change in velocity (we call this acceleration).

OpenStudy (faiqraees):

Yes I am aware that it is a derivative of dy/dx with respect to x. But how should I explain in Enlglish terms

OpenStudy (loser66):

@zepdrix thank you ;)

zepdrix (zepdrix):

derivative measures the movement* over time (we call this velocity), blah typo

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zepdrix derivative measures change in y with respect to x, so, second derivative measures change in (change in y with respect to x). It's measuring the change in the change. It makes a little more sense if you think of it in terms of movement I suppose. Like if you're measuring the movement of a particle over time... derivative measures the moment over time (we call this velocity), second derivative measures the change in velocity (we call this acceleration). \(\color{#0cbb34}{\text{End of Quote}}\) Yes that's exactly what I was thinking change in change in y with respect to x. But what about \[ \frac{dy^2}{dx^2 } \]

OpenStudy (faiqraees):

Change in change in y divided by change in change in x, perhaps?

zepdrix (zepdrix):

You're asking what the difference is between \(\large\rm \frac{d^2y}{dx^2}\) and \(\large\rm \frac{dy^2}{dx^2}\) ?

OpenStudy (faiqraees):

Yes

zepdrix (zepdrix):

This \(\large\rm \frac{dy^2}{dx^2}\) actually doesn't mean anything. It's not properly written. \(\large\rm \frac{d^2y}{dx^2}\) is the second derivative of y. \(\large\rm \frac{d^2y^2}{dx^2}\) is the second derivative of y^2. but if your top d doesn't have a second power, then it's just... not a thing D: It can't be second derivative in the bottom, and first derivative notation in the numerator. At least that's my understanding of it :d

zepdrix (zepdrix):

I guess you could do something like this,\[\large\rm \frac{dy^2}{dx^2}=\frac{dy^2}{dx}\cdot\frac{1}{dx}\]So it's the derivative of y^2 with respect to x, divided by the instantaneous rate of change of x. I don't know if that has any real meaning though.. hmm

OpenStudy (faiqraees):

\[\large\rm x^2=y^2\ \] \[\large\rm 2x ~ dx=2y~ dy\ \] \[\large\rm 2 ~ dx^2=2~ dy^2\ \] \[\large\rm 2 =2~ \frac{dy^2}{dx^2} \ \]

OpenStudy (faiqraees):

So according to you that dy^2/dx^2 doesn't mean anything what does it mean here?

OpenStudy (faiqraees):

And one last thing Is this true? And if not what does it equal to? \[\large\rm (\frac{dy}{dx})^2 =\frac{d^2y^2}{d^2x^2} \]

OpenStudy (loser66):

\(x^2=y^2\\x =\pm y\)

OpenStudy (sainath):

Wow people do math this early ._.

zepdrix (zepdrix):

Boy these are good questions :D I'm a little stumped lol. Still thinking...

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Loser66 \(x^2=y^2\\x =\pm y\) \(\color{#0cbb34}{\text{End of Quote}}\) I didn't take a root, I differentiated them \(\color{#0cbb34}{\text{Originally Posted by}}\) @Sainath Wow people do math this early ._. \(\color{#0cbb34}{\text{End of Quote}}\) Its 6 pm in my region \(\color{#0cbb34}{\text{Originally Posted by}}\) @zepdrix Boy these are good questions I'm a little stumped lol. Still thinking... \(\color{#0cbb34}{\text{End of Quote}}\) And I was thinking that you guys aren't replying because you consider my question s stupid. Thank you for clarification.

OpenStudy (loser66):

First off, your derivative is nonsense to me. if \(x^2=y^2\) \(\dfrac{d}{dx}(x^2)=\dfrac{d}{dy}(y^2)\) That gives us 2x = 2y It doesn't give 2xdx =2ydy

OpenStudy (loser66):

if you take second derivative, again, you get 2=2 , period

zepdrix (zepdrix):

You don't remember doing a little bit of that back in calculus Lose? :o Finding rate of change in terms of differentials.\[\large\rm y=3x+2\qquad\to\qquad dy=3dx\]

OpenStudy (faiqraees):

Oh apologies. Here is a modified version \[\large\rm x=y\\] \[\large\rm d/dx x=d/dyy\\] \[\large\rm d/dx dx=d/dy dy\\] \[\large\rm dx^2= dy^2\\] \[\large\rm 1 = \frac{dy^2}{dx^2} \\]

OpenStudy (faiqraees):

@Loser66 Your last comments was for the previous version right?

zepdrix (zepdrix):

I like your previous example better actually. \(\large\rm x^2=y^2\) Let's do it the normal way first. Differentiating each side `implicitly` with respect to x, \(\large\rm 2x=2yy'\) Solving for our derivative gives us,\[\large\rm y'=\frac{x}{y}\]squaring each side gives us,\[\large\rm (y')^2=\frac{x^2}{y^2}\] Let's check out what happens when we do it the other way,\[\large\rm 2xdx=2ydy\]Dividing each side by 2, squaring each side,\[\large\rm x^2dx^2=y^2dy^2\]Solving for ... something...\[\large\rm \frac{x^2}{y^2}=\frac{dy^2}{dx^2}\]So `in this particular case`, we determined that,\[\large\rm \frac{dy^2}{dx^2}=(y')^2=\left(\frac{dy}{dx}\right)^2\]

zepdrix (zepdrix):

So yesss it actually does have some meaning, ok my mistake. It's the square of the differentials... so when you divide them, it's the square of the derivative!

zepdrix (zepdrix):

Which does not necessarily have anything to do with the second derivative.

OpenStudy (faiqraees):

@zepdrix Thank you very much for explaining

zepdrix (zepdrix):

\[\large\rm \frac{dy^2}{dx^2}=\left[\frac{dy}{dx}\right]^2=\left[f'(x)\right]^2\]

zepdrix (zepdrix):

Which is not the same as this\[\large\rm \frac{d^2y}{dx^2}\]except in some rare cases.

OpenStudy (loser66):

To me, if y^2 =x^2 then dy^2/dx^2 =1

OpenStudy (loser66):

and d^2y/dx^2 =0

zepdrix (zepdrix):

Lose, remember doing differential math in integral calc with your u-subs? :)\[\large\rm u=x^2+3\qquad\to\qquad du=2x~dx\]

zepdrix (zepdrix):

I'm not sure of the exact meaning :P I just remember doing that sort of thing lol

OpenStudy (loser66):

let me explain!!

OpenStudy (loser66):

if y^2 =x^2 if you take derivative of y w.r.t x, then you get what you said above but if you take derivative of (y^2) with respect to the whole (x^2) then it is =1

OpenStudy (loser66):

after then, the second derivative of (y^2) w.r.t (x^2) =0 because it is derivative of a constant.

OpenStudy (faiqraees):

@zepdrix In your example, it sounds like this \[\large\rm u = x^2+3\ \] \[\large\rm \frac{d}{dx}u=\frac{d}{dx} (x^2+3)\ \] \[\large\rm \frac{du}{dx}=2x\ \] \[\large\rm du=2x dx\ \]

OpenStudy (faiqraees):

@zepdrix Thank you very much for explaining the meaning of \(\large\rm \frac{dy^2}{dx^2} \) and also how to incorporate LaTex so the terms come in line with text

zepdrix (zepdrix):

Ya this was a fun question :) I need to brush up on some Calc stuff